40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 289 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  2441   Mon Dec 21 19:24:29 2009 ranaUpdateASSOAF Model update and build instructions

I fit the MC1 -> MCL TF using vectfit4.m (from mDV). The wrapper file is mDV/extra/C1/ fitMC12MCL.m.

Plotted here are the data (RED), the fit (BLUE), and the residual x10 (GREEN).

For the magnitude plot, residual is defined as ------   res = 1 - fit / data

For the phase plot the residual is defined as -------    res =  phase(data)-phase(fit)

You can see that the agreement is very good. The phase match is better than 5 deg everywhere below 10 Hz.

This TF is so smooth that we could have probably done without using this, but its good to excercise the method anyway.

Attachment 1: mc12mcl.png
mc12mcl.png
  2440   Mon Dec 21 10:09:06 2009 jenneUpdateASSOAF Model update and build instructions
OAF stands for Online Adaptive Filtering. We use the same computer which was once the ASS. One of these days, we'd like to completely be rid of all things which refer to ASS, and make even the computer's name OAF.
  2439   Mon Dec 21 08:16:54 2009 steveMetaphysicsPEMjackhammering next door

At ITMX, on the CES side, 5 Ft  from the wall the jackhammer is on. The susses are holding well.

  2438   Mon Dec 21 07:30:58 2009 ???UpdateASSOAF Model update and build instructions

 What does OAF stand for? The entry doesn't say that. Also the acronym is not in the abbreviation page of the wiki.

Can anyone please explain that?

  2437   Mon Dec 21 02:22:31 2009 rana, jenne, kiwamuUpdateASSOAF Model update and build instructions

This allowed measuring the MC1 -> MCL TF finally. Its mostly flat. Data saved as Templates/OAF/OAF-MCLTF.xml

Attachment 1: Untitled.png
Untitled.png
  2436   Mon Dec 21 01:14:08 2009 ranaSummaryElectronicsNoise measurement of the Rai Weiss FET preamp box

 I shorted the input to the box and then put its output into the SR560 (low noise, G = 100, AC). I put the output of the SR560 into the SR785.

*** BTW, the 2nd channel of the SR785 is kind of broken. Its too noisy by a factor of 100. Needs to go back for repair once we get started in the vac.

The attached PNG shows its input-referred noise with the short.

The picture shows the inside of the box before I did anything. The TO-5 package metal can is the meaty super dual-FET that gives this thing all of its low noise power.

Picture_2.pngRWnoise.png

In the spectra on the right are two traces. The BLUE one is the noise of the box as I found it. The BLACK one is the noise after I replaced R1, R6, R7, & R10 with metal film resistors.

The offset at the output of the box with either an open or shorted input is +265 mV.

I think we probably should also replace R2, R3, & R1, but we don't have any metal film resistors lower than 100 Ohms in the kit...but hopefully Steve will read this elog and do the right thing.

Attachment 1: IMG_0242.JPG
IMG_0242.JPG
  2435   Sun Dec 20 23:42:44 2009 JenneUpdateIOONew Input Mode Matching Telescope

I've got most of the new Mode Matching Telescope figured out.  The scripts and an example result are at: MMT09 wiki  (Rather, the scripts are in the svn: MMT svn)

Issues still to be resolved: 

* We're getting pretty iffy 'angles' between tilt and translation when using the mode matching mirrors for steering.

* I haven't taken into account the astigmatism which occurs when you tilt the mode matching mirrors. 

The nifty thing about these scripts is that they take a look at the mode matching overlap:  For each possible mode matching solution it adds noise to all of the distances and radii of curvature during ~10,000 iterations and plots a histogram of the overlap so that we can see which solutions have a better chance of giving us the optimal overlap, even if we place the optics in slightly the wrong place.

 

I'd like to update the overlap part of the script with the astigmatism business:  do we lose goodness of overlap if we tilt the mirrors by a bit?  I think this will require redoing the overlap part with the X and Y directions separate.  Koji has done this in the past.  My current code assumes that the beam is always symmetric in X and Y. 

  2434   Sun Dec 20 21:39:40 2009 rana, jenne, kiwamuUpdateASSOAF Model update and build instructions

After a lot of headache, I got the OAF working again - read on for details.....................

Sometime last week, Jenne, Kiwamu, and I tried to update the OAF model to include the IIR "feed-around" path.

This path is in parallel to the existing FIR-based adaptive FXLMS stuff that Matt put in earlier. The reason for the

new path is that we want to try emulating the same FF technology which has been successful lately at LLO.

 

However, we were unable to make the ASS work after this work. Mostly, the build stuff worked fine, but we couldn't get DTT

to make a transfer function. The excitation channels could be selected and the excitation would actually start and get all the

way into MC1, but DTT would just hang on the first swep-sine measurement with no time-out error. Clearly our ASS building

documentation is no good. We tried using the instructions that Koji gave us for AAA, but that didn't completely work.

 

In particular, the 'make-uninstall-daq-ass' command gave this command:

[controls@c1ass advLigo]$ make uninstall-daq-ass
grep: target/assepics/assepics*.cmd: No such file or directory
Please make ass first
make: *** [uninstall-daq-ass] Error 1

re-arranging the order to do 'make-ass' first fixes this issue and so I have fixed this in the OAF Wiki.

 


The there's the whole issue with the tpchn_C3.par file. This contains all the test point definitions for the ASS/OAF machine. The main

IFO numbers are all in tpchn_C1.par and the OMC is all in tpchn_C2.par. When we do the usual build, in the 'make install-daq-ass' part:

[controls@c1ass advLigo]$ make install-daq-ass
Installing GDS node 3 configuration file
/cvs/cds/caltech/target/gds/param/tpchn_C3.par
Updating DAQ configuration file
/cvs/cds/caltech/chans/daq/C1ASS.ini

we get this .par file installed in the target area. The ACTUAL param file seems to actually be in /cvs/cds/caltech/gds/param  !!

 


Of course, it still doesn't work. That's because the standard build likes to point to /cvs/cds/caltech/gds/bin/awgtpman and the one that runs on

linux is actually /opt/gds/awgtpman. So I've now made a file called startup_ass.cmd.good which runs the correct one. However, the default build

will try to start the wrong one and we have to fix the 'startass' script to point to the correct one on each build. Running the correct awgtpman

allows us to get the TP data using tools like tdsdata, so far no luck with DTT.\

 


UPDATE (23:33): It turns out that it was my old nemesis, NTPD. c1ass had a /etc/ntp.conf file that was pointing at an ntp server called rana113. I

am not an NTP server; I don't even know what time it is. I have fixed the ntp.conf file by making it the same ass c1omc (it now points to nodus). After

this I set the date and time manually (sudo date -s "20 DEC 2009 23:27:45") and then restarted NTPD. It should now be fine even when

we reboot c1ass.

 

After all of this nonsense, I am able to get TP data from c1ass and take transfer functions between it and the rest of the world !

  2433   Sun Dec 20 14:34:24 2009 KojiUpdateSUSETMY watchdog tripped Sunday 5:00AM local

It seemed that the ETMY watchdog tripped early Sunday morning.
The reason is not known. I just looked at ETMX, but it seemed fine.

I called the control room just in case someone is working on the IFO.
Also I did not see any elog entry to indicate on going work there.

So, I decided to reset the watchdog for ETMY. And it is working fine again.

Attachment 1: Y.png
Y.png
  2432   Sat Dec 19 14:33:25 2009 KojiConfigurationComputersPDFlib lite / gnuplot 4.2.6 on Rosalba/Allegra

In order to enable 'set terminal pdf' in gnuplot on Rosalba/Allegra, I installed PDFlib lite and gnuplot v4.2.6. to them.
(PDFlib lite is required to build the pdf-available version of gnuplot)


Installation of PDFlib lite:

  • Building has been done at rosalba
  • Download the latest distribution of PDFlib lite from http://www.pdflib.com/products/pdflib-7-family/pdflib-lite/
  • Expand the archive. Go into the expanded directory
          tar zxvf PDFlib-Lite-7.0.4p4.tar.gz
          cd ./PDFlib-Lite-7.0.4p4
  • configure & make
          ./configure
          make
  • install the files to the system / configure the dinamic linker
          sudo make install
          sudo ldconfig

Installation of gnuplot:

  • Building has been done at rosalba
  • Download the latest distribution of gnuplot form http://www.gnuplot.info/
  • Expand the archive. Go into the expanded directory
          tar zxvf gnuplot-4.2.6.tar.gz
          cd ./gnuplot-4.2.6.tar.gz
  • configure & make
          ./configure --prefix=/cvs/cds/caltech/apps/linux/gnuplot
          make
          make install
  • Create symbolic links of the executable at
          /cvs/cds/caltech/apps/linux/bin
          /cvs/cds/caltech/apps/linux64/bin
  • Note: Although the original (non-PDF) gnuplot is still at
          /usr/bin/gnuplot
    new one is active because of the path setting
          rosalba:linux>which gnuplot
          /cvs/cds/caltech/apps/linux64/bin/gnuplot

 

  2431   Fri Dec 18 15:40:33 2009 KojiUpdateIOOMC2 spot centered / MCT QPD issue

This afternoon I felt like saying hello to the input mode cleaner. So I decided to center the spot on MC2.

Motivation

MC has 6 alignment dofs. 4 of them are controlled by the WFSs. Remaining 2 appears at the spot position on MC2.
If the spot on the MC2 is fixed, the beam hits the same places of three mirrors. If the mirrors are completely fixed
in terms of the incident beam, I suppose the reflected beam is also fixed. This makes the WFS spots more stable.
Then I feel better.

Today's goal is to confirm the behaviour of MC such as dithering amplitude, response of the couplings to the alignment,
behavior of the WFS, and the transmitted power.


Method

1) Turned off MC auto locker. Turned off MC WFS as the WFS servos disturbs my work.
2) Dithered MC2 in Pitch and Yaw using DTT. There looks elliptic filter (fc=28Hz) in the ASC path, I used 20Hz-ish excitations.
- C1:SUS-MC2_ASCPIT_EXC 100cnt_pk@19Hz
- C1:SUS-MC2_ASCYAW_EXC 100cnt_pk@22Hz
3) Looked at C1:SUS-MC2_MCL_OUT to find the peaks at 19Hz and 22Hz. These are caused by alignment-length coupling.
If they are minimized I assume the spot is somehow centered on MC2.
Note: This may not be the true center. The suspension response should be investigated. But this is a certain reporoducible spot position.
Note: I should use ezcademod in order to obtain the phase information of the dither result.

4) Move MC2 Pitch for certain amount (0.01cnt) by the alignment slider. Align MC1/MC3 to have max transmittion.
5) If the Pitch peak got lower, the direction of 4) was right. Go further.
5') If the Pitch peak got higher, the direction of 4) was wrong. Go the other direction.

6) Repeat 4)&5) for Yaw.


Result

After the adjustment, the couplings got lower about 10 times. (Sorry! The explanation is not so scientific.)
Next time I (or someone) should make a script to do it and evaluate the coupling by the estimated distance of the spot from the center of the mirror (the center of the rotation).
I have not seen visible change in the spectrum of C1:SUS-MC2_MLC_OUT.


MCT QPD issue

By the spot centering, I could expected to see some improvement of the transmittion. But in reality, there was no change.
In fact, the transmittion power was getting down for those weeks.

I checked WFS and MCT paths. Eventually I found that a couple of possible problems:
1) MCT Total output varies more than 10% depending on the spot position on MCT QPD.
2) Just before the QPD, there is a ND1 filter.
This may suggest that:
a) Four elemtns of the MCT QPD have different responses.
b) The ND filter is causing a fringe.

So far I aligned the ND filter to face the beam. The reflection from the filter was blocked at a farther place.
Still the output varies on the spot position. Probably I have to look at the QPD someday.

So far the spot on the QPD was defined so that I get the maximum output from the QPD. This is about 8.8.
As I touched the steering mirrors, the X and Y outputs of the QPD are no longer any reference.

For now, I closed the PSL table. The full IFO was aligned.

  2430   Thu Dec 17 23:27:23 2009 ranaUpdatePEMRanger Noise: sim w. Rai FET box as readout

I have started measuring the low frequency noise of the FET front end + LT1128 low noise preamp from Rai Weiss. It has a very low input current noise because its FET based, which is not surprising. It is also a fairly low voltage noise box - the best measured ones have an input referred noise of ~0.35 nV/rHz.

Today I measured the noise of the one we have down to 0.1 Hz. It looks like a good candidate for a Geophone readout (e.g. Ranger or GS-13 or perhaps the L-4C). Because I didn't thermally shield any of this stuff, the broadband noise is ~0.8 nV/rHz. The low frequency corner is ~15 Hz.

I attach the LISO simulation of the voltage noise referred to the input. The circuit is described in this entry.

We can probably do better than this if we package it a little better or give it time to warm up or use metal film resistors inside. Even as it is, however, it would allow us to reach the thermal noise of the Ranger (or GS-13) down to 0.1 Hz.

This should be ~1.5 or 2x better than the LT1012 based readout at 1 Hz and 10x better down at 0.1 Hz (c.f. T0900457).

Attachment 1: ranger.pdf
ranger.pdf
  2429   Thu Dec 17 19:03:14 2009 AlbertoHowToComputersNodus sluggish

Quote:

The elog has been quite slow for the last two days. The cause is nodus, that has been slowing down the access to it.

I looked at the list of the running processes on nodus by typing the command prstat, which is the equivalent for Solaris of the Linux "top". I didn't see any particular process that might be absorbing too many resources.

I remember Rana fixing the same problem in the past but couldn't find his elog entry about that. Maybe we should just restart nodus, unless someone has some other suggestion.

 Problem solved. Nodus and the elog are running fine. It's just that the elog takes some time to make a preview of complex pdf attachments, like those in Kiwamu's entry 2405.

  2428   Thu Dec 17 17:13:50 2009 AlbertoOmnistructureEnvironmentSTACIS stuff
One of the electronics benches is currently occupied by the STACIS equipment.
We need that table If no one is working on the STACIS anymore, it should be removed from there.
  2427   Thu Dec 17 09:30:08 2009 AlbertoHowToComputersNodus sluggish

The elog has been quite slow for the last two days. The cause is nodus, that has been slowing down the access to it.

I looked at the list of the running processes on nodus by typing the command prstat, which is the equivalent for Solaris of the Linux "top". I didn't see any particular process that might be absorbing too many resources.

I remember Rana fixing the same problem in the past but couldn't find his elog entry about that. Maybe we should just restart nodus, unless someone has some other suggestion.

  2426   Thu Dec 17 07:47:29 2009 JenneUpdateWienerFilteringL1 DARM Static Wiener Filtered data

This surface plot is the same as the previous one, with a little more data than I had previously. 

This time around, I also include the "BLRMS" plots for this data.  The first one takes each residual and normalizes it by the DARM_CTRL signal at that time, separates the spectra into bands, and integrates underneath the spectra within that band.  The second one is the raw DARM_CTRL signal's spectra at each time, and integrates under the spectra for each band, and the third BLRMS plot does the same thing for the residuals.  Unfortunately, these plots don't have the same handy black stripe during time which I don't analyze that the spectrogram utilizes.

From the second BLRMS plot we can see that the large red splotch in the spectrogram is due to higher noise in the DARM spectrum, and that (by looking at the Ratio BLRMS plot) the Wiener filter still does a pretty good job during this time.  I expect that for later times when the seismic (or something) event is gone, the Wiener filter will continue performing almost as well as it had been initially.

Again, once the script finishes applying the filter to the many ten minute chunks (the huge time drain is the data fetching, so this shouldn't be a limiting factor for using Wiener filters online), I'll post a final plot.

Attachment 1: L1darmComp_17Dec2009_6day_residualsNormSurfacePlot.png
L1darmComp_17Dec2009_6day_residualsNormSurfacePlot.png
Attachment 2: L1darmComp_17Dec2009_6day_ratioBLRMS.png
L1darmComp_17Dec2009_6day_ratioBLRMS.png
Attachment 3: L1darmComp_17Dec2009_6day_rawBLRMS.png
L1darmComp_17Dec2009_6day_rawBLRMS.png
Attachment 4: L1darmComp_17Dec2009_6day_residualsBLRMS.png
L1darmComp_17Dec2009_6day_residualsBLRMS.png
  2425   Thu Dec 17 02:57:08 2009 JenneUpdateWienerFilteringL1 DARM Static Wiener Filtered data

This is perhaps best put in the LLO elog, but I'm not yet a 'person' there, so I can't write to their elog (yet another thing for the eternal to-do list).  So for now, we're putting things here...

This isn't totally finalized, but I do want to get what I have posted before I hop on a plane in the morning.  Mostly it just needs more time to run, to make the plot longer.  Hopefully I'll be able to edit this in the morning and have a longer-duration plot. 

What's plotted:

This spectrogram shows the amplitude spectra of L1:LSC-DARM_CTRL, after being subtracted via a Static Wiener Filter.  Each spectra is normalized by the very first one, which was created from the same data that was used to determine the Wiener Filter.  The X-axis is time.  The Y-axis is frequency, and the Color/Z-axis is amplitude in dB.  I'm only looking at Science Mode time, so other times when the IFO isn't in science mode, I plot a black stripe to fill in the plot.  The start time of the plot is 83675598, which is Jul 08 2006 06:33:04 UTC. 

Why?

The idea is to see that the filter does equally well a long time after it was created, as when it was initially made.  This will help tell us how often it is useful to recompute the Wiener filters.  Less often is nice, because redoing the Wiener filters may also include remeasuring the high precision transfer functions...if the filter isn't working as well anymore it may be because the transfer function has changed ever so slightly. 

How the plot is created / the background story:

I use one hour of DARM_CTRL data and the following seismometer channels to create an optimal Wiener Filter (pem indicates L0:PEM- , sei indicates L1:SEI- , and lsc indicates L1:LSC- ) :

chans = {[pem 'EX_SEISX'],...
         [pem 'EX_SEISY'],...
         [pem 'EX_SEISZ'],...
         [pem 'EY_SEISX'],...
         [pem 'EY_SEISY'],...
         [pem 'EY_SEISZ'],...
         [pem 'LVEA_SEISX'],...
         [pem 'LVEA_SEISY'],...
         [pem 'LVEA_SEISZ'],...
         [sei 'LVEA_STS2_X'],...
         [sei 'LVEA_STS2_Y'],...
         [sei 'LVEA_STS2_Z'],...
         [sei 'ETMX_STS2_X'],...
         [sei 'ETMX_STS2_Y'],...
         [sei 'ETMX_STS2_Z'],...
         [sei 'ETMY_STS2_X'],...
         [sei 'ETMY_STS2_Y'],...
         [sei 'ETMY_STS2_Z'],...
         [lsc 'DARM_CTRL']};

I then apply this one filter to ten minute chunks of science mode data, for some long period of time.  The game plan is to have a month long plot, but it takes a while to fetch all of the data in separate 10min intervals (~45sec per iteration, times ~3000 iterations), so this plot isn't a full month.  Even if I don't get a chance to plot a full month by Thursday morning, it'll go up here within the next few days. The particular times chosen have the most science mode data within a 30 day period.  I can easily run the code for some other time, if there is a known time (or season) which might be more interesting.  For the spectrogram plot, I then normalize each amplitude spectra by the first one, which comes from the first ten minutes in the hour which was used to make the filter.  This makes it easier to see how the filter's efficacy changes over time.

The analogous analysis for Hanford is in the 40m elog: 1606.  The Hanford stuff in the elog has some cool BLRMS plots also, but I'm not sure that they're so helpful when I only have a few days of L1 data so far.  I'll do those and add them later.

Conclusions:

I can't really say anything yet about the long-term efficacy of a Wiener Filter for LLO yet, since my code hasn't finished filtering my one month of S5 L1 data.  It definitely looks like (so far) that there was a big seismic event around the (arbitrarily defined) 'Day 4'. 

Attachment 1: L1darmCompPlot_17Dec2009_4daysLong.png
L1darmCompPlot_17Dec2009_4daysLong.png
  2424   Wed Dec 16 20:29:08 2009 ranaUpdateCOCETM Coating study

This plot shows the Transmission for 532 and 1064 nm as a function of the thickness of the SiO2 layer.

i.e. the thickness is constrained so that the optical thickness of the SiO2 and Ta2O5 pair is always 1/2 of a wavelength.

The top layer of the mirror is also fixed in this plot to be 1/2 wave.

This plot shows the result for 17 pairs. For 16 pairs, we are unable to get as low as 15 ppm for the 1064 nm transmission.

Attachment 1: layerfrac.png
layerfrac.png
  2423   Wed Dec 16 11:55:47 2009 ranaUpdateABSLUniversal PDH Box Servo Filters

 

 To me, they both look stable. I guess that the phase has to go to -180 deg to be unstable.

Why does the magnitude go flat at high frequencies? That doesn't seem like 1/f.

How about a diagram of what inputs and outputs are being measured and what the gain knob and boost switch settings are?

  2422   Wed Dec 16 11:46:25 2009 AlbertoUpdateABSLAbsl PLL Open Loop Gain

Yesterday I measured the Open Loop Gain of the PLL in the absolute length experiment. The servo I used was that of the old Universal PDH box.

The OLG looks like this:

OldBoxOLTF.png

The UGF is at 10 KHz.

  2421   Wed Dec 16 11:21:20 2009 AlbertoUpdateABSLUniversal PDH Box Servo Filters
Yesterday I measured the shape of the servo filter of both the old and the new Universal PDH boxes.
Here they are compared.

NewandOlfFilterTF.png

The way the filter's transfer function has been measured is by a swept sine between the "SERVO INPUT" and the "PIEZO DRIVE OUTPUT" connection on the box front panel. The spectrum analyzer used for the measurement is the SR785 and the source amplitude is set at 0.1V.

The two transfer functions are clearly different. In particular the old one looks like a simple integrator, whereas the new one already includes some sort of boost.

That probably explains why the new one is unable to lock the PLL. Indeed what the PLL needs, at least to acquire lock, is an 1/f filter.

I thought the two boxes were almost identical, at least in the filter shapes. Also the two schematics available in the DCC coincide.

Attachment 1: NewandOlfFilterTF.png
NewandOlfFilterTF.png
  2420   Tue Dec 15 21:39:34 2009 AlbertoUpdateABSLbrief summary of this afternoon's measurements
I took measurements of the open loop gain of the AbsL PLL with the old Universal PDH Box.
I Also measured the filter shape of both the new and the old PDH box.
I'm going to plot the results in a nice form tomorrow morning.
For who's interested, the PLL UGF was at 10KHz.
 
I can't lock the PLL with the new PDH box. Measuring its filter's shape, as suggested by Koji, I found out that it differs from the old one. That despite the fact that the two boxes should share the same circuit schematic. O,r at least, that is what it looks like from the schematics in the DCC.
I need to understand whether that is intentional and, if that was the case, what kind of use  Rich Abbott designed it for.
 
Tomorrow I'm going to post in the elog the filter's transfer functions too.
 
Before leaving the lab I closed the auxiliary laser's mechanical shutter.
  2419   Tue Dec 15 17:16:22 2009 KojiUpdateGeneralTable distance measurements

During the vent we have tried to measure the distances of the optical tables for BS-ITMX and BS-ITMY.
We need to take into account the difference between the AutoCAD drawing and the reality.

X direction distance of the table center for BS and ITMX:
84.086" (= 2135.8mm)
(This is 84.0000" in AutoCAD drawing)

Y direction distance of the table center for BS and ITMX:
83.9685" (= 2132.8mm)
(This is 83.5397" in AutoCAD drawing)

We used two scales attached each other in order to measure the distance of the certain holes on the tables.

We got more numbers that were estimated from several separated measurements.
I think they were not so accurate, but just as a record, I also put the figure as an attachment 2.

Attachment 1: Table_distance_by_metal_scale.pdf
Table_distance_by_metal_scale.pdf
Attachment 2: Table_distance_by_chambers.pdf
Table_distance_by_chambers.pdf
  2418   Tue Dec 15 05:29:31 2009 AlbertoUpdateGeneralArm Cavity Poles measured again after cleaning the optics last week

 

 The Y arm cavity pole moved down by 130 Hz, whereas the X arm moved by only 34 Hz. I wonder if that is because Kiwamu spent much more time on cleaning ITMY than on any other optic.

  2417   Tue Dec 15 00:02:10 2009 ranaUpdatePEMNoise of the Ranger SS-1 Seismometer

I wanted to see what the noise of the Ranger seismometer should be. I used LISO and file ranger.fil (in our LISO SVN) to calculate the voltage noise referred to the input. In this model, we represent the EMF from the moving magnet in the coil as a voltage source at 'nin' which drives the coil impedance. This is the same approach that Brian Lantz uses in his noise modeling of the GS-13 (PDF is on our Ranger wiki page).

In the simulation, I used the OP27 as a placeholder for the SR560 that we use (I don't know the current noise of the SR560). To do this, I use the new 'inputnoise' feature in LISO (its in the README, but not in the manual).

You can see that we would be limited by the input current noise of the OP27. So we would do a lot better if we used an FET based readout amp like the AD743 (or equivalent) or even better using the new multi-FET readout circuit that Rich Abbott has developed. Clearly, its also silly to have a load resistance in there - I put it in because the manual says to do it, but all it does is damp the mass and reduce the size of the signal.

#  Noise sim for the Ranger SS-1 seismometer
#
#                            \
#                           | \            
#           n2- - - ----  - |  \          
#            |    |         | op1>-- n4 - r4 -- no
#           Rg   RL     n3- |  /     |           
#       n1 - |    |     |   | /      |                
#           Lg    |     |    /       |                        
#            |    |     | - - - R2 - -                    
#           nin  gnd   R1 
#                       |
#                      gnd

We previously measured the Ranger's self noise by locking it down.

The 1/f^3 noise that we see below 1 Hz is roughly consistent with the noise model: to get from my plot into meters you have to multiply by:

(1 + f)^2

----------

340 * f^2

P.S. Secret PDF handshake: You can make your non-compliant applications like LISO or DTT produce a thumbnailing PDF by using Acrobat to open the file and export it as PDF/A.

In the second attachment, I have used an OPA827 (new low-noise FET input amp from TI) as the readout amplifier. This seems like a good choice - main drawback is that Digikey backordered my OPA827s by 19 weeks!

Attachment 1: rangerx.pdf
rangerx.pdf
Attachment 2: ranger827.pdf
ranger827.pdf
  2416   Mon Dec 14 22:32:56 2009 KojiUpdateGeneralArm cavity loss ~ result

I like to ask someone to review the calculation on the wiki.

In the calculation, the round trip loss and the front mirror T are the unknown variables.
The end mirror T of 10ppm was assumed. (End mirror T)+(Round trip loss) is almost invariant, and T_end does not change the other results much.

Arm cavity loss measurement (Dec. 14, 2009)

X Arm:
  • Arm visibility (given): 0.897 +/- 0.005 (20 pts) (2.5%UP!!)
  • Cut off freq (given): 1616 +/- 14 [Hz] (2.1%UP!!)
  • Finesse (derived): 1206 +/- 10 (2.1%UP!!)
  • Round Trip loss (estimated): 127 +/- 6 [ppm] (28%DOWN!!)
  • Front Mirror T (estimated): 0.00506 +/- 0.00004
Y Arm:
  • Arm visibility (given): 0.893 +/- 0.004 (20 pts) (2.1%UP!!)
  • Cut off freq (given): 1590 +/- 4 [Hz] (8.2%UP!!)
  • Finesse (derived): 1220 +/- 3 (8.2%UP!!)
  • Round Trip loss (estimated): 131 +/- 6 [ppm] (37%DOWN!!)
  • Front Mirror T (estimated): 0.00500 +/- 0.00001 

Previous measurement (Oct 07, 2009 & Nov 10, 2009)

X Arm:  

  • Arm visibility (given): 0.875 +/- 0.005 (34 pts)
  • Cut off freq (given): 1650 +/- 70 [Hz]
     
  • Finesse (derived): 1181 +/- 50
  • Round Trip loss (estimated): 162 +/- 10 [ppm]
  • Front Mirror T (estimated): 0.0051 +/- 0.0002

Y Arm: 

  • Arm visibility (given): 0.869 +/- 0.006 (26 pts)
  • Cut off freq (given): 1720 +/- 70 [Hz]
     
  • Finesse (derived): 1128 +/- 46
  • Round Trip loss (estimated): 179 +/- 12 [ppm]
  • Front Mirror T (estimated): 0.0054 +/- 0.0002
  2415   Mon Dec 14 19:33:04 2009 AlbertoUpdateGeneralArm Cavity Poles measured again after cleaning the optics last week

Last week we vented and we cleaned the main optics of the arm cavities.

I measured the frequency of the cavity poles for both the arm cavities to see how they changed (see previous elog entry 2226). These the results:

fp_X = 1616 +/- 14 Hz

fp_Y = 1590 +/- 4 Hz

The error is the statistical error that I got with the Matlab NonLinearLeastSquare fitting function.
 
I calculated the cavity pole frequencies by measuring the transfer function between a photodiode located at the end of the arms (either X or Y) and another photodiode placed after the mode cleaner. Both diodes where Thorlabs PDA255.
(Last time, I had measured that the pair of diode had a flat calibration).
 
With the SR785 I measured the transfer function by exciting the OMC_ISS_EXC input cable.
For both arms I set to 1V the excitation amplitude. I repeated the measurements for different excitation amplitudes without observing any changes.
I then fitted the data with the NonLinearLeastSquare function of matlab. The script I wrote to do that is attached to this entry in a compressed file.
The files also contains the PDF with the output plots and the data from both set of measurements performed before and after the cleaning.
The data is commented in a file called measurements.log.
 
In the end I disabled again the test switch on the ISS MEDM screen.
I disconnected the excitation cable from the OMC_ISS_EXC input cable.
I removed the photodiode that measured the Mode Cleaner transmission from the PSL clearing the way for the beam to get back to its path to the RFAM photodiode.
Attachment 1: XarmTF01_fit.pdf
XarmTF01_fit.pdf
Attachment 2: YarmTF01_fit.pdf
YarmTF01_fit.pdf
Attachment 3: ArmCavityFinesseMeasurment.tar.gz
  2414   Mon Dec 14 15:18:18 2009 JenneUpdatelorearmLoss script ran....results confidential

I ran the armLoss script for both Xarm and Yarm.  The results are confidential, pending the completion of Alberto's cavity pole/finesse measurement due to the 'bet' as to what the new losses are after the drag wiping.

If you're the kind of person who likes to look at their Chrismas presents, the log files with the results are in the usual place for this script: /scripts/LSC/loss-ARM-GPStime.log  (loss-Y-944865071.log and loss-X-944865946.log)

  2413   Mon Dec 14 14:16:14 2009 AlbertoUpdateTreasureLOCKSTARS

Quote:

Good job guys. What I did was saying "I don't know", "Maybe", and "Ants...".

Now you can proceed to measurements for the visibility and the cavity pole! 

Quote:

[Jenne, Kiwamu, Koji]

We got the IFO back up and running!  After all of our aligning, we even managed to get both arms locked simultaneously.  

 I'm going to do it right now.

  2412   Mon Dec 14 13:17:33 2009 robUpdateTreasureWe are *ROCKSTARS* ! IFO is back up

 

 

Attachment 1: two-thumbs-up.jpeg
two-thumbs-up.jpeg
  2411   Mon Dec 14 13:08:33 2009 KojiUpdateTreasureLOCKSTARS

Good job guys. What I did was saying "I don't know", "Maybe", and "Ants...".

Now you can proceed to measurements for the visibility and the cavity pole! 

Quote:

[Jenne, Kiwamu, Koji]

We got the IFO back up and running!  After all of our aligning, we even managed to get both arms locked simultaneously.  

  2410   Mon Dec 14 12:13:52 2009 JenneUpdateTreasureWe are *ROCKSTARS* ! IFO is back up

[Jenne, Kiwamu, Koji]

We got the IFO back up and running!  After all of our aligning, we even managed to get both arms locked simultaneously.  Basically, we are awesome. 

 This morning, we did the following:

*  Turned on the PZT High voltages for both the steering mirrors and the OMC.  (For the steering mirrors, turn on the power, then hit "close loop" on each.  For the OMC, hit Output ON/OFF).

*  Looked at the PZT strain gauges, to confirm that the PZTs came back to where they had been.  (Look at the snapshot of C1ASC_PZT_Al)

*  Locked all components of the PSL (This had already been done.)

*  Removed beam dump which was blocking the PSL, and opened the PSL mechanical shutter.  Light into the IFO!

*  Locked the Mode Cleaner.  The auto-locker handled this with no problem.

*  Confirm that light is going through the Faraday.  (Look at the TV sitting on top of MC13 tank...it shows the Faraday, and we're hitting the input of the Faraday pretty much dead-on).

*  Look at IP_ANG and IP_POS.  Adjust the steering mirrors slightly to zero the X&Y readings on IP_ANG.  This did not change the PZTs by very much, so that's good.

*  Align all of the Core Optics to their OpLev positions.

*  On the IFO_Align screen, save these positions.

*  Run the IFO_Configure scripts, in the usual order.  (Xarm, Yarm, PRM, DRM).  Save the appropriate optics' positions after running the alignment scripts.  We ended up running each alignment script twice, because there was some residual misalignment after the first iteration, which we could see in the signal as viewed on DataViewer (Either TRX, TRY, or SPOB, for those respective DoFs).

*  Restore Full IFO.

*  Watch the beauty of both arms and the central cavity snapping together all by themselves!  In the attached screenshot, notice that TRX and TRY are both ~0.5, and SPOB and AS166Q are high.  Yay!

Conclusions: 

*  The wiping may have helped.  While aligning X and Y separately, TRX got as high as ~1.08, and TRY got as high as 0.98  This seems to be a little bit higher than it was previously.

*  Since everything locked up in pretty short order, and the free swinging spectra (as measured by Kiwamu in elog 2405) looks good, we didn't break anything while we were in the chambers last week.  Excellent.

*  We are now ready for a finesse measurement to tell us more quantitatively how we did with the wiping last week.

 

Attachment 1: Jenne14Dec09_IFOlocked.png
Jenne14Dec09_IFOlocked.png
  2409   Mon Dec 14 11:21:23 2009 steveOmnistructureEnvironmentAnts in the coffee maker

 

 We still had some ants visiting the sink area this morning. These ants seem to be addicted to our our Peet's coffe

Spectracide: Bug Stop insect killer was sprayed. Please wash your eating dishes well ! and keep area clean.

 

  2408   Mon Dec 14 00:37:28 2009 KojiOmnistructureEnvironmentAnts in the coffee maker

I made a short stop at the 40m on Sunday night and found that hundreds ants are in the coffee maker.
I removed ants around the sink and washed the coffee maker.

It looked the ants were everywhere in the lab tonight. They seemed to prefer warm places like in the coffee maker and below the coffee mill.
So, I recommend that Steve should confirm there is no ants in the coffee maker again before the first coffee of the week is made.
Othewise they will add some more acidity to your cup.

 

  2407   Sun Dec 13 23:18:09 2009 ranaSummaryIOODisplacement noise on the PSL table

For the Laser Gyro, I wondered how much mechanical noise we might get with a non-suspended cavity. My guess is that the PMC is better than we could do with a large ring and that the MZ is much worse than we could do.

Below 5 Hz, I think the MZ is "wind noise" limited. Above 10 Hz, its just ADC noise in the readout of the PZT voltage.

Attachment 1: mz.pdf
mz.pdf
  2406   Sun Dec 13 20:50:45 2009 ranaSummaryIOOMach Zender Calibration

I ramped the MZ PZT (with the loop disabled on the input switch) to calibrate it. Since the transmission has been blocked, I used the so-called "REFL" port of the MZ to do this.

The dark-to-dark distance for the MZ corresponds to 2 consecutive destructive interferences. Therefore, that's 2 pi in phase or 1 full wavelength of length change in the arm with the moving mirror.

Eyeballing it on the DTT plot (after lowpassing at 0.1 Hz) and using its cursors, I find that the dark-to-dark distance corresponds to 47.4 +/- 5 seconds.

So the calibration of the MZ PZT is 88 +/- 8 Volts/micron.

Inversely, that's a mean of 12 nm / V.

why am I calibrating the MZ? Maybe because Rob may want it later, but mainly because Koji won't let me lock the IFO.

Apparently, we haven't had a fast channel for any of the MZ board. So I have temporarily hooked it up to MC_DRUM at 21:13 and also turned down the HEPA. Now, let's see how stable the MZ and PMC really are overnight.

EDIT: it railed the +/- 2V ADCwe have so I put in a 1:4 attenuator via Pomona box. The calibration of MC_DRUM in terms of MZ_PZT volts is 31.8 cts/V.

So the calibration of MC_DRUM1 in meters is: 0.38 nm / count


Attachment 1: Untitled.png
Untitled.png
  2405   Sun Dec 13 17:43:10 2009 kiwamuUpdateSUSfree swinging spectra (vacuum)

The free swinging spectra of ITMs, ETMs, BS, PRM and SRM were measured under the vacuum-condition. The attachment are measured spectra.

It looks there are nothing wrong because no significant difference appear from the past data and the current data (under atmosperic pressure).

So everything is going well.

Attachment 1: summary_FreeSwinging_vacuum.pdf
summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf
  2403   Sat Dec 12 07:36:56 2009 ranaHowToElectronicsHow to Measure the Length of a Cable: Interferometry

Need to measure the length of the cable, but too lazy to use a measuring tape?

Then you too can become an expert cable length measurer by just using an RF signal generator and a scope:

  1. Disconnect or short (not 50 Ohm term) the far side of the cable.
  2. Put a T on the near side of the cable.
  3. Drive the input of the T with your signal source.
  4. Look at the output of the T with the scope while sweeping the signal source's frequency knob.

The T is kind of acting like a beamsplitter in an asymmetric length Michelson in this case. Just as we can use the RF phase shift between the arms to measure the Schnupp asymmetry, we can also use a T to measure the cable length. The speed of light in the cable is documented in the cable catalog, but in most cases its just 66% of the speed of light in the vacuum.

  2402   Fri Dec 11 19:51:13 2009 AlbertoConfigurationLSC166 LO Disconnected

Quote:

 

 Seems like very strange cable loss numbers. The Heliax is lossier than the RG-174? I wonder how these compare with the specs in the cable catalog?

In my last entry there was a typo for the measurement of the Heliax at 55 MHz. I corrected it. It was 1.084 dB instead of 1.084 dB/m.
For the Heliax I don't have the measurement of the loss per meter since I don't know the cable actual length.
 
Except for that, I checked the values I found on the Internet.
My measurements for the RG-174 seem comparable to the loss specified in the catalog (here): 6.6dB in 100ft @ 55 MHz, that is 0.22 dB/m, that compare with 0.155 dB/m that I measured.

I did the measurement on a 4.33 meter long cable with SMA connectors at the ends.

  2401   Fri Dec 11 17:36:37 2009 kiwamuUpdateGeneralIFO restoring plan

Alberto, Jenne, Kiwamu

 

We together will lead the IFO restoring and the following is our plan.

- - - - -|- - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -

#.0     |  measuring the free swinging spectra                     (weekend by kiwamu)   DONE

#.1     |  turn ON the PZTs for steering mirror and so on.         (Dec.14 Mon.) DONE

#. 1            |    lock around PSL  DONE

#.2     |  deal with mechanical shutter                            (Dec.14 Mon.)DONE

#.3     |  lock MCs                                                (Dec.14 Mon.)DONE

#.4     |  align the IFO                                           (Dec.15 Tue.)DONE

#.5     |  lock full IFO                                           (Dec.15 Tue.)DONE

- - - - -|- - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -

 

Thank you.

  2400   Fri Dec 11 15:21:40 2009 steveConfigurationVACpumpdown#67 is completed

Quote:

Oplev positions before and after drag wiped arm TMs as of yesterday. Slow-mode pumpdown has started with 3/4 turn opened  RV1 valve at 8am today.

 Pump down is completed. Valve configuration is VACUUM NORMAL. CC1 pressure is in the ~8 e-5 torr   PSL output shutter is opened.

Attachment 1: pd7h.jpg
pd7h.jpg
Attachment 2: oplevpd8h.jpg
oplevpd8h.jpg
  2399   Fri Dec 11 14:19:22 2009 KojiConfigurationVACpumpdown has started

Wait, Wait, Wait. You are moving too fast. Go one by one.

Check the PZTs, the MC, initial pointings, IFO mirrors, some of the partial locks, and maybe some momentary full locks?
Once the recover of the IFO is declared, you can proceed to the measurements.

I hope the grad students can take this precious opportunity to have their fun time for restoring everything by themselves.

Quote:

 I'm leaving the lab now for less than 2 hours. I should be back in time for when the pumping is finished so that I can measure the finesse again.

 

  2398   Fri Dec 11 14:12:32 2009 ranaConfigurationLSC166 LO Disconnected

 

 Seems like very strange cable loss numbers. The Heliax is lossier than the RG-174? I wonder how these compare with the specs in the cable catalog?

  2397   Fri Dec 11 13:18:17 2009 AlbertoConfigurationVACpumpdown has started

Quote:

Oplev positions before and after drag wiped arm TMs as of yesterday. Slow-mode pumpdown has started with 3/4 turn opened  RV1 valve at 8am today.

 I'm leaving the lab now for less than 2 hours. I should be back in time for when the pumping is finished so that I can measure the finesse again.

  2396   Fri Dec 11 11:42:26 2009 AlbertoConfigurationLSC166 LO Disconnected

Quote:

They must not be dBm, must be dB

Quote:

Quote:

I temporarily disconnected the Heliax cable that brings the 166MHz LO to the LSC rack.

I'm doing a couple of measurement and I'll put it back in as soon as I'm done.

 These are the losses I measured on a RG-174 cable for the two frequencies that we're planning to use in the Upgrade:

@11MHz Loss=0.22dBm/meter

@55MHz Loss=0.41dBm/meter

(The cable was 2.07m long. The input signal was +10dBm and the output voltages at the oscilloscope where: Vpk-pk(11MHz)=1.90V, Vpk-pk(11MHz)=1.82V )

 

I apologize for the lack of correctness on the units in yesterday's elog entry, but I wasn't very sharp last night.

I repeated the measurement today, this time also making sure that I had a 50ohm input impedance set in the scope. These the results for the losses.

RG-174 Cable 0.053 dB/m @ 11MHz  0.155 dB/m @ 55 MHz

 I also measured the losses in the Heliax cable going from the 166 MHz LO to the LSC rack:

166MHz LO Heliax 0.378 dB @ 11MHz  1.084 dB @ 55 MHz

 

  2395   Fri Dec 11 09:30:09 2009 KojiConfigurationLSC166 LO Disconnected

They must not be dBm, must be dB

Quote:

Quote:

I temporarily disconnected the Heliax cable that brings the 166MHz LO to the LSC rack.

I'm doing a couple of measurement and I'll put it back in as soon as I'm done.

 These are the losses I measured on a RG-174 cable for the two frequencies that we're planning to use in the Upgrade:

@11MHz Loss=0.22dBm/meter

@55MHz Loss=0.41dBm/meter

(The cable was 2.07m long. The input signal was +10dBm and the output voltages at the oscilloscope where: Vpk-pk(11MHz)=1.90V, Vpk-pk(11MHz)=1.82V )

 

  2394   Fri Dec 11 08:35:04 2009 steveConfigurationVACpumpdown has started

Oplev positions before and after drag wiped arm TMs as of yesterday. Slow-mode pumpdown has started with 3/4 turn opened  RV1 valve at 8am today.

Attachment 1: wiping.jpg
wiping.jpg
  2393   Thu Dec 10 18:31:44 2009 AlbertoConfigurationLSC166 LO Disconnected

Quote:

I temporarily disconnected the Heliax cable that brings the 166MHz LO to the LSC rack.

I'm doing a couple of measurement and I'll put it back in as soon as I'm done.

 These are the losses I measured on a RG-174 cable for the two frequencies that we're planning to use in the Upgrade:

@11MHz Loss=0.22dBm/meter

@55MHz Loss=0.41dBm/meter

(The cable was 2.07m long. The input signal was +10dBm and the output voltages at the oscilloscope where: Vpk-pk(11MHz)=1.90V, Vpk-pk(11MHz)=1.82V )

  2392   Thu Dec 10 18:27:48 2009 MottUpdateGeneralUpdated R & T Measurements

Attached are updated plots of the T&R Measurements for a variety of mirrors, and diagrams for the setup used to make the measurements.

T is plotted for the 1064 nm measurement, since these mirrors are highly reflective at 1064, and either R or R&T are plotted for the 532 nm measurement, depending on how larger the R signal is.

As with the previous set of plots, the error bars here are purely statistical, and there are certainly other sources of error not accounted for in these plots.  In general, the T measurement was quite stable, and the additional errors
are probably not enormous, perhaps a few percent.

The mirrors are:

Y1-1037-00.50CC

Y1-2037-45S

Y1-2037-45P

Y1S-1025-0

Y1S-1025-45

 

Attachment 1: Y1S-0.png
Y1S-0.png
Attachment 2: Y1S-45.png
Y1S-45.png
Attachment 3: Y45P.png
Y45P.png
Attachment 4: Y45S.png
Y45S.png
Attachment 5: Y150CC.png
Y150CC.png
Attachment 6: Setup.png
Setup.png
  2391   Thu Dec 10 17:13:36 2009 kiwamuUpdateSUSFree swiging spectra under the atmospheric pressure

The free swinging spectra of ITMs, ETMs, BS, PRM and SRM were measured last night in order to make sure that nothing wrong have happened by the wiping.

I think there are nothing wrong with ITMs, ETMs, BS, PRM and SRM successfully.

For the comparison, Yoichi's figure in his elog entry of Aug.7 2008 is good, but in his figure somehow PRM spectrum doesn't look correct.

Anyway, compared with his past data, there are no significant changes in the spectra. For PRM which has no counterpart to compare with, its shape of spectra looks similar to any other spectra. So I think PRM is also OK. The measured spectra are attached below.

Indeed the current data are still showing smaller peak height for their pitch and yaw modes (roughly factor of 10 ).
 
Attachment 1: summary_freeswing.pdf
summary_freeswing.pdf summary_freeswing.pdf summary_freeswing.pdf summary_freeswing.pdf summary_freeswing.pdf summary_freeswing.pdf summary_freeswing.pdf
ELOG V3.1.3-